# Face Forgery Detection with Elaborate Backbone

### v1

@@ 4.1.2 Implementation Details
Following MoCo v3, MAE, and BEiT v2 settings, we pre-train ViT on our real-face dataset using $4$x A100 GPUs until full convergence. We use RetinaFace [121] for face cropping, resize the cropped faces to $224$×$224$, and sample $8$ frames at equal intervals from each video. Fine-tuning is performed on $2$x GTX 3090 GPUs with a batch size $64$, using the Lamb optimizer [122] and a learning rate $55e-$ 5 with cosine decay for $60$ epochs. Notably, backbones pre-trained on real faces with more robust facial feature representation capabilities converge in just $10$ epochs.
@@ 4.1.3 Evaluation Metrics and Protocols
Evaluation Metrics. Following many existing FFD and PAD works, we use the Area Under the Receiver Operating Characteristic Curve (AUC) and Half Total Error Rate (HTER) as evaluation metrics, respectively. Video-level scores are computed by averaging frame-level scores for comparison.
@@ 4.2 Comparison of FFD with Pre-trained Backbones
We conduct experiments to evaluate FFD performance of backbones with various learning approaches, network architectures, and datasets. The quantitative evaluation results are shown in Table I. We use the backbone with randomly initialized network parameters as a baseline (Unpre-trained) to highlight the importance of pre-training for FFD task.
Unpre-trained Backbones. We utilize unpre-trained backbones with different architectures, including ResNet, Xception, EfficientNet, and ViT, and fine-tune them on FF++ dataset. Their poor AUC results, averaging around $61$%, underscore the significance of pre-trained backbones in FFD.
Learning Approaches. We employ mainstream backbones pre-trained with SL and SSL on ImageNet-1K (i.e., SL-1K and SSL-1K) to evaluate the impact of learning approaches in FFD. As shown in Table I, SL-1K achieves substantially better FFD performance than unpre-trained backbones, demonstrating the indispensability of pre-trained backbones. Furthermore, we use ResNet-50 and ViT-B, commonly used in SSL, as the backbone network architectures. The comparison results in 3.1 and 3.2 of Table I indicate that SSL gradually surpasses SL in FFD performance with the improvement of pre-training methods, such as the momentum encoder/contrast in CL equipped in SimSiam, SimCLR, and MoCo, and the MIM pretext task equipped in MAE, BEiT, and ConMIM. This trend is consistent with their performance on downstream ImageNet tasks, proving that appropriate learning approaches are also crucial for FFD.
TABLE I: Quantitative comparison of FFD with different backbones. The average AUC (Avg) across three cross-datasets is provided for reference. 1K and Face refer to ImageNet-1K and our real-face dataset, respectively. 1K&Face denotes a backbone pre-trained on 1K and then further pre-trained on Face. Bold, bold, and italics indicate the best overall, second-best, and best in each section, respectively.
@@ 3.1 SSL-1K (ResNet-50)
[…]
@@ 3.2 SSL-1K (ViT-B)
[…]
@@ 3.3 SSL-Face (ViT-B)
MoCo v3 [91] $98.41$ $86.30$ $78.51$ 83.38  $82.73$
MAE [24] $98.22$ $85.43$ $78.33$ $80.02$   $81.26$
BEiT v2 [25] 98.93 88.62 80.99 $81.71$ $\bm{83.77}$
@@ 3.4 SSL-1K&Face (ViT-B)
MoCo v3 [91] $98.65$ $84.98$ $79.13$ $\bm{85.72}$  $83.28$
MAE [24] $98.37$ $85.85$ $78.04$ $78.07$  $80.65$
BEiT v2 [25] $\bm{99.14}$ $\bm{89.18}$ $\bm{84.02}$ $\bm{86.48}$ $\bm{86.56}$
Network Architectures. We then evaluate the effect of different network architectures, which serve as the backbone’s infrastructure. The comparison results in SSL-1K of Table I show that the size of the network parameters under the same architecture has little impact on FFD performance. Meanwhile, ViT substantially outperforms ResNet under SSL, e.g., MoCo v3 with ViT-B surpasses ResNet-50 by $3.77$% on average AUC across three cross-datasets. Thus, an excellent network architecture combined with an appropriate learning approach will significantly promote the backbone’s baseline performance in FFD.
@@
Extensive empirical studies on backbones demonstrate that embedding both general and face-specific knowledge into the network through pre-training methods aligned with FFD challenges is essential for advancing generalization.
@@ 4.3 Comparison with State-of-the-arts
We move on to our competitive backbone fine-tuning framework, which enhances the backbone’s forgery detection capabilities through the competitive learning mechanism. We compare our model’s generalization and adaptability with state-of-the-art FFD methods, such as FTCN [72], SBI [55], RECCE [69], CORE [33], UIA-ViT [21], and LSDA [20]. Furthermore, we use BEiT v2 pre-trained on the ImageNet-1K and our real-face dataset as the baseline (BEiT v2’) and introduce SBI to verify its data augmentation performance.
@@ 4.3.1 Generalization Evaluation
We first evaluate the generalization capability of various FFD methods on cross-datasets: CDF, DFDC, and FFIW. We also present the average AUC across these three datasets for a clear comparison. Table II shows the quantitative comparison results, as can be seen: (1) previous FFD methods exhibit varying and unsatisfactory performance on cross-datasets with diverse forgery patterns, indicating that generalization to unseen cross-datasets remains a significant challenge in FFD task; (2) except SBI [55], which uses sophisticated augmentation strategies instead of FF++ forged videos during training, even our baseline BEiT v2’ significantly outperforms existing methods in generalization, including FTCN [72] and AltFreezing [57], which utilize additional temporal information. These results highlight the significance of a proficient backbone in FFD models, as well as the indispensable role of analyzing the correspondent relationship between the capabilities of different backbones and FFD problems; (3) Ours and Ours+SBI achieve state-of-the-art generalization performance, especially on the more diverse and complex forgery patterns in DFDC and FFIW, which present greater challenges and differ significantly from the within-dataset FF++. These results demonstrate that our competitive backbone fine-tuning framework effectively equips the backbone with a powerful ability to uncover implicit forgery cues across different face synthesis algorithms. Additionally, compared to UIA-ViT [21], which also uses ViT-B as the network architecture, our framework achieves a $10.81$% performance improvement, demonstrating our method’s superior ability to unlock ViT’s potential.
@@
CDF DFDC FFIW Avg
FTCN [72] CVPR 2021 $87.88$ $71.31$ $87.62$ $82.27$
SBI* [55] CVPR 2022 $\underline{92.87}$ $71.96$ $83.22$ $82.68$
RECCE [69] CVPR 2022 $71.58$ $64.53$ $73.51$ $69.87$
CORE [33] CVPR 2022 $80.10$ $70.53$ $78.42$ $76.35$
UIA-ViT [21] ECCV 2022 $83.75$ $74.34$ $75.82$ $77.97$
D-adv [124] TIP 2022 $82.77$ $73.91$ $72.84$ $76.51$
UCF [17] ICCV 2023 $77.45$ $70.68$ $66.16$ $71.43$
CADDM [18] CVPR 2023 $84.39$ $70.46$ $85.72$ $80.19$
AltFreezing [57] CVPR 2023 $89.23$ $72.38$ $75.79$ $79.13$
LSDA* [20] CVPR 2024 $91.10$ $77.00$- --
Choi et al.* [19] CVPR 2024 $89.00$- -- --
BEiT v2’ - $89.18$ 84.0284.02{84.02}84.02 86.4886.48{86.48}86.48 $86.56$
Ours - $90.46$ $\underline{84.90}$ $\underline{90.97}$ $\underline{88.78}$
Ours+SBI - 95.35 85.45 91.82 90.87
Additionally, we further evaluate the generalization of different FFD models on our $9$ cross-datasets. The quantitative comparison in Table III shows that both our baseline BEiT v2’ and Ours outperform other FFD methods, highlighting the importance of a backbone with real face knowledge and the effectiveness of competitive forgery cue extraction. However, all listed FFD methods in Table III exhibit poor generalization performance on StyleGAN3-generated faces, mainly due to unseen entire face forgeries during training and the high realism of these fakes. Therefore, we employ a small number of StyleGAN3-generated forged faces ($20$ images) to fine-tune different methods further, enabling verification of their rapid adaptability to unseen forgery techniques. The bottom row of Table III presents the adaptability performance, where RECCE [69] and BEiT v2’ show greater improvement by modeling real face features, while Ours demonstrates even more impressive performance with a $21.45$% improvement, proving its superior practicality.
@@
Forged Faces RECCE CORE D-adv CADDM BEiT v2 BEiT v2’ Ours
@@ 4. Face Reenactment
Avg AUC (%) $66.82$ $53.21$ $56.94$ $60.86$ $72.63$ $\underline{81.90}$ 83.86
Forged Faces RECCE CORE D-adv CADDM BEiT v2 BEiT v2’ Ours
@@ 4.3.2 Adaptation Evaluation
TABLE IV: Quantitative adaptation comparison of different FFD models by further fine-tuning on a few images ($10$-Shot, $20$-Shot) from unseen datasets. These models have been fine-tuned on the source dataset (FF++). bold means the best, and bold means the second-best. $\uparrow$ and $\downarrow$ indicate an increase and decrease relative to the baseline (gray), respectively. Higher values on the diagonal represent greater performance and adaptation.
@@
($10$/$20$-Shot) FF++ CDF DFDC FFIW Avg FF++ CDF DFDC FFIW Avg
RECCE [69] FF++ ${\color[rgb]{.5,.5,.5}97.91}$ ${\color[rgb]{.5,.5,.5}71.58}$ ${\color[rgb]{.5,.5,.5}64.53}$ ${\color[rgb]{.5,.5,.5}73.51}$ ${\color[rgb]{.5,.5,.5}76.88}$ ${\color[rgb]{.5,.5,.5}97.91}$ ${\color[rgb]{.5,.5,.5}71.58}$ ${\color[rgb]{.5,.5,.5}64.53}$ ${\color[rgb]{.5,.5,.5}73.51}$  ${\color[rgb]{.5,.5,.5}76.88}$
CDF $89.25\downarrow$ $67.78\downarrow$ $65.45\uparrow$ $66.59\downarrow$ $72.27\downarrow$ $93.58\downarrow$ $70.45\downarrow$ $61.95\downarrow$ $67.50\downarrow$ $73.37\downarrow$
@@
FFIW $97.66\downarrow$ $67.73\downarrow$ $65.22\uparrow$ $73.96\uparrow$ $76.14\downarrow$ $94.62\downarrow$ $70.31\downarrow$ $66.55\uparrow$ $77.49\uparrow$ $77.24\uparrow$
CORE [33] FF++ ${\color[rgb]{.5,.5,.5}97.66}$ ${\color[rgb]{.5,.5,.5}80.10}$ ${\color[rgb]{.5,.5,.5}70.53}$ ${\color[rgb]{.5,.5,.5}78.42}$ ${\color[rgb]{.5,.5,.5}81.68}$ ${\color[rgb]{.5,.5,.5}97.66}$ ${\color[rgb]{.5,.5,.5}80.10}$ ${\color[rgb]{.5,.5,.5}70.53}$ ${\color[rgb]{.5,.5,.5}78.42}$ $\ {\color[rgb]{.5,.5,.5}81.68}$
@@
D-adv [124] FF++ ${\color[rgb]{.5,.5,.5}98.31}$ ${\color[rgb]{.5,.5,.5}82.77}$ ${\color[rgb]{.5,.5,.5}73.91}$ ${\color[rgb]{.5,.5,.5}72.84}$ ${\color[rgb]{.5,.5,.5}81.96}$ ${\color[rgb]{.5,.5,.5}98.31}$ ${\color[rgb]{.5,.5,.5}82.77}$ ${\color[rgb]{.5,.5,.5}73.91}$ ${\color[rgb]{.5,.5,.5}72.84}$ $\ {\color[rgb]{.5,.5,.5}81.96}$
@@
BEiT v2 FF++ ${\color[rgb]{.5,.5,.5}99.16}$ ${\color[rgb]{.5,.5,.5}89.48}$ ${\color[rgb]{.5,.5,.5}81.00}$ ${\color[rgb]{.5,.5,.5}77.61}$ ${\color[rgb]{.5,.5,.5}86.81}$ $\bm{{\color[rgb]{.5,.5,.5}99.16}}$ ${\color[rgb]{.5,.5,.5}89.48}$ ${\color[rgb]{.5,.5,.5}81.00}$ ${\color[rgb]{.5,.5,.5}77.61}$ ${\color[rgb]{.5,.5,.5}86.81}$
@@
BEiT v2’ FF++ ${\color[rgb]{.5,.5,.5}99.14}$ ${\color[rgb]{.5,.5,.5}89.13}$ ${\color[rgb]{.5,.5,.5}84.02}$ ${\color[rgb]{.5,.5,.5}86.48}$ ${\color[rgb]{.5,.5,.5}89.71}$ ${\color[rgb]{.5,.5,.5}99.14}$ ${\color[rgb]{.5,.5,.5}89.13}$ ${\color[rgb]{.5,.5,.5}84.02}$ ${\color[rgb]{.5,.5,.5}86.48}$ ${\color[rgb]{.5,.5,.5}89.71}$
@@
Ours FF++ $\bm{{\color[rgb]{.5,.5,.5}99.36}}$ ${\color[rgb]{.5,.5,.5}90.46}$ ${\color[rgb]{.5,.5,.5}84.90}$ ${\color[rgb]{.5,.5,.5}90.97}$ ${\color[rgb]{.5,.5,.5}91.42}$ $\bm{{\color[rgb]{.5,.5,.5}99.36}}$ ${\color[rgb]{.5,.5,.5}90.46}$ ${\color[rgb]{.5,.5,.5}84.90}$ ${\color[rgb]{.5,.5,.5}90.97}$ ${\color[rgb]{.5,.5,.5}91.42}$
CDF $\bm{99.25}\downarrow$ $\bm{91.33}\uparrow$ $86.17\uparrow$ $91.74\uparrow$ $92.12\uparrow$ $99.10\downarrow$ $\bm{91.31}\uparrow$ $85.48\uparrow$ $91.82\uparrow$ $91.93\uparrow$

### v2

@@ 4.1.2 Implementation Details
Following MoCo v3, MAE, and BEiT v2 settings, we pre-train ViT on our real-face dataset using 44x A100 GPUs until full convergence. We use RetinaFace [26] for face cropping, resize the cropped faces to 224224×$224$, and sample $8$ frames at equal intervals from each video. Fine-tuning is performed on 22x GTX 3090 GPUs with batch size $64$, using the Lamb optimizer [130] and a learning rate 5e55e-5 with cosine decay for $60$ epochs. Notably, backbones pre-trained on real faces with more robust facial feature representation capabilities converge in just $10$ epochs. In comparisons, we use $64$ frames per video for temporal-based FFD methods because they require more frames, and $8$ frames per video for spatial-based FFD methods, which are independent of video length.
@@ 4.1.3 Evaluation Metrics and Protocols
Evaluation Metrics. Similar to existing FFD and PAD works, we use the Area Under the Receiver Operating Characteristic Curve (AUC) and Half Total Error Rate (HTER) as evaluation metrics, respectively. Our work presents video-level AUC results, where frame prediction probabilities are averaged within each video to obtain its final probability, followed by AUC calculation over the test set.
@@ 4.2 Comparison of FFD with Pre-trained Backbones
We evaluate the importance of pre-trained backbones for FFD models by analyzing three core pre-training configurations: learning approaches, network architectures, and datasets. Table I presents the quantitative results, with the unpre-trained backbone as the baseline (Unpre-trained).
Unpre-trained Backbones. We utilize unpre-trained backbones with different architectures, including ResNet, Xception, EfficientNet, and ViT, and fine-tune them on FF++ dataset. These models yield an average AUC of around 6161%, indicating predictions close to random guessing.
TABLE I: Quantitative comparison of FFD with different backbones. The average AUC (Avg.) across three cross-datasets is provided for reference. 1K and Face refer to ImageNet-1K and our real-face dataset, respectively. 1K&Face denotes a backbone pre-trained on 1K and then further pre-trained on Face. Bold, bold, and italics indicate the best overall, second-best, and best in each section, respectively.
@@ 3.1 SSL-1K (ResNet-50)
Supervised $\color[rgb]{.5,.5,.5}97.95$}   $\color[rgb]{.5,.5,.5}81.24$} $\color[rgb]{.5,.5,.5}72.23$} $\color[rgb]{.5,.5,.5}74.21$} $\color[rgb]{.5,.5,.5}75.89$}
@@ 3.2 SSL-1K (ViT-B)
Supervised $\color[rgb]{.5,.5,.5}97.09$}   $\color[rgb]{.5,.5,.5}83.39$} $\color[rgb]{.5,.5,.5}76.77$} $\color[rgb]{.5,.5,.5}71.17$} $\color[rgb]{.5,.5,.5}77.11$}
@@
3.3 VLM and STM (ViT-B)
CLIP [85] $98.87$   $\bm{89.48}\$78.86 $77.43$ 81.92
BLIP [60] $98.60$   $87.90$ $78.21$ 77.82 $81.31$
SigLIP [132] $98.71$   $84.07$ $78.34$ $75.69$ $79.37$
VideoSwin [70] $\bm{99.37}$   $80.62$ $77.58$ $76.74$ $78.31$
VideoMAE [102] $99.32$   $84.43$ $74.24$ $77.41$ $78.69$
UMT [68] $\bm{99.43}\$  $85.22$ $74.91$ $76.15$ $78.76$
Uniformer v2 [61] $99.19$   $84.42$ $75.94$ $75.72$ $78.69$
@@ 3.4 SSL-Face (ViT-B)
MoCo_Face $98.41$   $86.30$ $78.51$ 83.38 $82.73$
MAE_Face $98.22$   $85.43$ $78.33$ $80.02$ $81.26$
BEiT_Face 98.93   88.62 80.99 $81.71$ $\bm{83.77}$
@@ 3.5 SSL-1K&Face (ViT-B)
FaceMoCo $98.65$   $84.98$ $79.13$ $\bm{85.72}$ $83.28$
FaceMAE $98.37$   $85.85$ $78.04$ $78.07$ $80.65$
FaceBEiT (FB) 99.14   $\bm{89.18}$ $\bm{84.02}\$$\bm{86.48}\$$\bm{86.56}\$
Learning Approaches. We employ mainstream backbones pre-trained with SL and SSL on ImageNet-1K (i.e., SL-1K and SSL-1K) to evaluate the impact of learning approaches in FFD. As shown in Table I, SL-1K achieves over a 1010% performance improvement compared to unpre-trained backbones, demonstrating the indispensability of pre-trained backbones with better image feature representation. Besides, as shown by ViT-B (384) in SL-1K, increasing the input image resolution from $224$ to $384$ benefits FFD performance but also increases computational cost.
Moreover, we employ ResNet-50 and ViT-B as backbone architectures to evaluate the impact of pre-training with different SSL methods. The comparison results in 3.1 and 3.2 of Table I indicate that SSL gradually surpasses SL in FFD performance with the improvement of pre-training methods, such as the momentum encoder/contrast in CL equipped in SimSiam, SimCLR, and MoCo, and the MIM pretext task equipped in MAE, BEiT, and ConMIM. The results of SSL-1K highlight the effectiveness of introducing SSL pre-training backbones and the necessity of exploring image feature representation approaches tailored for FFD.
Network Architectures. We then evaluate the effect of different network architectures, which serve as the backbone’s infrastructure. The comparison results in SL-1K of Table I show that the size of the network parameters under the same architecture has little impact on FFD performance. Meanwhile, ViT significantly outperforms ResNet in SSL-1K, e.g., MoCo v3 with ViT-B surpasses ResNet-50 by 3.773.77% on average AUC across three cross-datasets. Thus, an excellent network architecture combined with an appropriate learning approach will significantly promote the backbone’s baseline performance in FFD.
Pre-training Datasets. We construct a large-scale real-face dataset to pre-train backbones, aiming to enhance their ability to represent facial features by alleviating the gap between nature images and facial images. We choose pre-training methods of MoCo v3, MAE, and BEiT v2, which are prominent in CL and MIM for single-modal pre-training and also perform excellently in FFD. We did not use BEiT v3 because its pre-training requires an image-text dataset.
In addition, we evaluate vision–language backbones (CLIP [85], BLIP [60], SigLIP [132]) pre-trained on larger-scale datasets such as WIT [85] and WebLI [132], as well as spatio-temporal backbones (VideoSwin [70], VideoMAE [102], UMT [68], Uniformer v2 [61]) with supervised pre-training on Kinetics [54].
@@
Extensive empirical studies on backbones demonstrate that embedding both general and face-specific knowledge through SSL pre-training methods is crucial for enhancing facial representation capabilities, thereby improving the fundamental generalization in FFD models.
@@ 4.3 Comparison with State-of-the-arts
We move on to our competitive backbone fine-tuning framework, which enhances the backbone’s forgery detection capabilities through the competitive learning mechanism. We compare our model’s generalization and adaptability with state-of-the-art FFD methods, such as FTCN [137], SBI [95], RECCE [5], CORE [80], UIA-ViT [140], LSDA [121], Choi et al[19], LLA+SBI [79], Yan et al[126], DFD-FCG [43], M2F2-Det [38], and TFCU [40]. Furthermore, our framework fine-tunes FB and FB++ to obtain Competitive FaceBEiT (CFB) and Competitive FaceBEiT++ (CFB++), respectively. FB is pre-trained on ImageNet-1K and our real-face dataset, whereas FB++ extends FB with additional real-face pre-training. We also introduce SBI to verify its data-augmentation performance.
@@ 4.3.1 Generalization Evaluation
We first evaluate the generalizability of various FFD methods on $5$ classic cross-datasets and present the average AUC across these datasets for a clear comparison. Table II shows the quantitative results, as can be seen: (1) previous FFD methods exhibit varying and unsatisfactory performance on cross-datasets with diverse forgery patterns, indicating that generalization to unseen cross-datasets remains a significant challenge. (2) Both CFB and TFCU [40] utilize FB as the backbone and achieve relatively stable performance across $5$ common cross-datasets (all exceeding 8484%), significantly outperforming existing methods. These results highlight the importance of backbone facial representation capability for effective fine-tuning in FFD methods. It is worth noting that TFCU relies on video temporal information, while our models only requires spatial information from a single frame, enabling support for both image and video inputs. (3) CFB++ achieves further performance gains by leveraging FB++ pre-trained on a larger real-face dataset and achieves state-of-the-art performance when combined with the SBI data augmentation strategy (CFB++ $w/$ SBI), especially on the more diverse and complex forgery patterns in DFDC and FFIW, which pose greater challenges and differ significantly from the within-dataset FF++. These results demonstrate that our fine-tuning framework effectively equips the backbone with a powerful ability to uncover implicit forgery cues across different face synthesis algorithms. For more analytical experiments, see Section 4.5.
@@
CDF DFDC FFIW DFDCP DFD Avg.
[…]
CORE $^{22^{\prime}} 80$.1080.10 $70.53$ $78.42$ $72.19$ $92.05$ $78.66$
[…]
M2F2-Det $^{25^{\prime}} 83$.2283.22 $76.46$ $81.78$ $71.22$ $88.98$ $80.33$
[…]
CFB $90.46$ $84.90$ $90.97$ $87.16$ $\underline{96.21}$ $89.94$
CFB++ $90.99$ $\underline{86.79}$ $90.72$ $87.17$ 96.52 $90.44$
CFB w/ SBI $\underline{95.35}$ $85.45$ $\underline{91.82}$ $\underline{94.13}$ $96.04$ $\underline{92.56}$
CFB++ w/ SBI 95.46 88.66 92.31 94.42 $95.92$ 93.35
@@
Forged Faces    RECCE    CORE    D-adv    CADDM    BEiT v2    FB    CFB
@@ 4. Face Reenactment
Avg. AUC (%)  $66.82$  $53.21$  $56.94$  $60.86$  $72.63$  $\underline{81.90}$  83.86
Forged Faces    RECCE  CORE  D-adv  CADDM  BEiT v2  FB  CFB
@@
We evaluate different FFD models on the $9$ cross-datasets we created. Table III shows that both our baseline FB and CFB outperform other FFD methods, emphasizing the importance of a backbone with real face knowledge and the effectiveness of competitive forgery cue extraction. However, all listed FFD methods in Table III exhibit poor generalization performance on StyleGAN3-generated faces, mainly due to unseen entire face forgeries during training and the high realism of these fakes. Therefore, we employ a small number of StyleGAN3-generated forged faces ($20$ images) to fine-tune different methods further, enabling verification of their rapid adaptability to unseen forgery techniques. The bottom row of Table III presents the adaptability performance, where RECCE [5] and FB show greater improvement by modeling real face features, while CFB demonstrates even more impressive performance with a 21.4521.45% improvement, proving its superior practicality. Additionally, we also validate the superiority of our method on DF40 across various face forgery algorithms. As shown in Table IV, our method achieves state-of-the-art performance with an average AUC improvement of 2.912.91%.
TABLE IV: Quantitative generalization comparison of various FFD methods on the DF40 dataset with different face forgery algorithms [123]. “†” denotes the author’s trained model, and “‡”indicates re-implementation using public code.
Method SadTalker[133] FOMM[96] FaceDancer[90] MobileSwap[119] SimSwap[11] InSwapper[49] UniFace[117] Avg.
FTCN [137] $87.76$ $87.23$ $85.08$ $77.46$ $94.56$ $97.12$ $98.63$ $89.69$
SBI [95] $78.51$ $99.12$ $76.43$ $\underline{99.23}$ $95.54$ $89.54$ $93.73$ $90.30$
RECCE [5] $86.24$ $98.52$ $81.94$ $96.87$ $80.28$ $93.39$ $93.59$ $90.12$
CORE [80] $87.59$ $98.48$ $86.05$ $99.12$ $91.19$ $93.83$ $95.40$ $93.09$
UIA-ViT [140] $76.50$ $93.64$ $84.07$ $89.87$ $72.37$ $89.77$ $89.78$ $85.14$
D-adv [113] $86.19$ $98.88$ $80.23$ $95.44$ $81.62$ $88.36$ $93.47$ $89.17$
UCF [124] $71.27$ $85.10$ $83.87$ $90.47$ $61.64$ $80.72$ $80.18$ $79.04$
CADDM [30] $54.64$ $79.36$ $71.56$ $96.62$ $92.32$ $75.73$ $89.84$ $80.01$
AltFreezing [112] $88.28$ $76.47$ $87.75$ $87.66$ $82.84$ $88.88$ $86.92$ $85.54$
CFM [74] $83.87$ $98.66$ $92.65$ $98.63$ $89.37$ $93.60$ $96.40$ $93.31$
LLA+SBI  [79] $82.14$ $92.85$ $87.14$ $97.15$ $92.12$ $90.00$ $90.71$ $90.30$
M2F2-Det [38] $83.61$ $97.53$ $93.55$ $98.33$ $92.18$ $94.83$ $97.35$ $93.91$
TFCU [40] $\underline{88.34}$ $\underline{99.29}$ $94.66$ $98.84$ $94.99$ $96.61$ $98.77$ $95.93$
CFB $87.94$ $99.15$ 95.23 $99.18$ 95.92 97.22 98.93 $\underline{96.22}$
CFB++ 88.98 99.39 $\underline{94.93}$ 99.52 $\underline{95.76}$ $\underline{97.14}$ $\underline{98.91}$ 96.38
@@ 4.3.2 Adaptation Evaluation
TABLE V: Quantitative adaptation comparison of different FFD models by further fine-tuning on a few images ($10$-Shot, 2020-Shot) from unseen datasets. These models have been fine-tuned on the source dataset (FF++). bold means the best, and bold means the second-best. $\uparrow$ and $\downarrow$ indicate an increase and decrease relative to the baseline (gray), respectively. Higher values on the diagonal represent greater performance and adaptation.
@@
(1010/2020-Shot) FF++ CDF DFDC FFIW Avg. FF++ CDF DFDC FFIW Avg.
RECCE [5] FF++ $\color[rgb]{.5,.5,.5}97.91$}\ \ \ $\color[rgb]{.5,.5,.5}71.58$}\ \ \ $\color[rgb]{.5,.5,.5}64.53$}\ \ \ $\color[rgb]{.5,.5,.5}73.51$}\ \ \ $\color[rgb]{.5,.5,.5}76.88$}\ \ \ $\color[rgb]{.5,.5,.5}97.91$}\ \ \ $\color[rgb]{.5,.5,.5}71.58$}\ \ \ $\color[rgb]{.5,.5,.5}64.53$}\ \ \ $\color[rgb]{.5,.5,.5}73.51$}\ \ \ $\color[rgb]{.5,.5,.5}76.88$}\ \ \
CDF $89.25\downarrow$ $67.78\downarrow$ $5.45\uparrow 66$$6.59\downarrow$ $72.27\downarrow$ $93.58\downarrow$ $70.45\downarrow$ $61.95\downarrow$ $67.50\downarrow$ $73.37\downarrow$
@@
FFIW $97.66\downarrow$ $67.73\downarrow$ $65.22\uparrow$ $73.96\uparrow$ $76.14\downarrow$ $94.62\downarrow$ $70.31\downarrow$ $66.55\uparrow$ $7.49\uparrow 77$$7.24\uparrow$
CORE [80] FF++ $\color[rgb]{.5,.5,.5}97.66$}\ \ \ $\color[rgb]{.5,.5,.5}80.10$}\ \ \ $\color[rgb]{.5,.5,.5}70.53$}\ \ \ $\color[rgb]{.5,.5,.5}78.42$}\ \ \ $\color[rgb]{.5,.5,.5}81.68$}\ \ \ $\color[rgb]{.5,.5,.5}97.66$}\ \ \ $\color[rgb]{.5,.5,.5}80.10$}\ \ \ $\color[rgb]{.5,.5,.5}70.53$}\ \ \ $\color[rgb]{.5,.5,.5}78.42$}\ \ \ $\color[rgb]{.5,.5,.5}81.68$}\ \ \
@@
D-adv [113] FF++ $\color[rgb]{.5,.5,.5}98.31$}\ \ \ $\color[rgb]{.5,.5,.5}82.77$}\ \ \ $\color[rgb]{.5,.5,.5}73.91$}\ \ \ $\color[rgb]{.5,.5,.5}72.84$}\ \ \ $\color[rgb]{.5,.5,.5}81.96$}\ \ \ $\color[rgb]{.5,.5,.5}98.31$}\ \ \ $\color[rgb]{.5,.5,.5}82.77$}\ \ \ $\color[rgb]{.5,.5,.5}73.91$}\ \ \ $\color[rgb]{.5,.5,.5}72.84$}\ \ \ $\color[rgb]{.5,.5,.5}81.96$}\ \ \
@@
M2F2-Det [38] FF++ $\color[rgb]{.5,.5,.5}99.13$}\ \ \ $\color[rgb]{.5,.5,.5}83.22$}\ \ \ $\color[rgb]{.5,.5,.5}76.46$}\ \ \ $\color[rgb]{.5,.5,.5}81.78$}\ \ \ $\color[rgb]{.5,.5,.5}85.15$}\ \ \ $\bm{{\color[rgb]{.5,.5,.5}99.13}}\ \ \$$\color[rgb]{.5,.5,.5}83.22$}\ \ \ $\color[rgb]{.5,.5,.5}76.46$}\ \ \ $\color[rgb]{.5,.5,.5}81.78$}\ \ \ $\color[rgb]{.5,.5,.5}85.15$}\ \ \
CDF $98.26\downarrow$ $80.22\downarrow$ $70.64\downarrow$ $72.67\downarrow$ $80.40\downarrow$ $98.00\downarrow$ $78.69\downarrow$ $68.02\downarrow$ $71.05\downarrow$ $78.94\downarrow$
DFDC $96.91\downarrow$ $81.68\downarrow$ $2.54\downarrow 77$$7.71\downarrow$ $82.21\downarrow$ $96.52\downarrow$ $81.15\downarrow$ $70.94\downarrow$ $81.06\downarrow$ $82.42\downarrow$
FFIW $94.53\downarrow$ $78.98\downarrow$ $74.28\downarrow$ $89.66\uparrow$ $84.36\downarrow$ $92.93\downarrow$ $77.99\downarrow$ $74.02\downarrow$ $90.06\uparrow$ $83.75\downarrow$
BEiT v2 FF++ $\color[rgb]{.5,.5,.5}99.16$}\ \ \ $\color[rgb]{.5,.5,.5}89.48$}\ \ \ $\color[rgb]{.5,.5,.5}81.00$}\ \ \ $\color[rgb]{.5,.5,.5}77.61$}\ \ \ $\color[rgb]{.5,.5,.5}86.81$}\ \ \ $\bm{{\color[rgb]{.5,.5,.5}99.16}}\ \ \$$\color[rgb]{.5,.5,.5}89.48$}\ \ \ $\color[rgb]{.5,.5,.5}81.00$}\ \ \ $\color[rgb]{.5,.5,.5}77.61$}\ \ \ $\color[rgb]{.5,.5,.5}86.81$}\ \ \
@@
FB FF++ $\color[rgb]{.5,.5,.5}99.14$}\ \ \ $\color[rgb]{.5,.5,.5}89.13$}\ \ \ $\color[rgb]{.5,.5,.5}84.02$}\ \ \ $\color[rgb]{.5,.5,.5}86.48$}\ \ \ $\color[rgb]{.5,.5,.5}89.71$}\ \ \ $\color[rgb]{.5,.5,.5}99.14$}\ \ \ $\color[rgb]{.5,.5,.5}89.13$}\ \ \ $\color[rgb]{.5,.5,.5}84.02$}\ \ \ $\color[rgb]{.5,.5,.5}86.48$}\ \ \ $\color[rgb]{.5,.5,.5}89.71$}\ \ \
@@
CFB FF++ $\bm{{\color[rgb]{.5,.5,.5}99.36}}\ \ \$$\color[rgb]{.5,.5,.5}90.46$}\ \ \ $\color[rgb]{.5,.5,.5}84.90$}\ \ \ $\color[rgb]{.5,.5,.5}90.97$}\ \ \ $\color[rgb]{.5,.5,.5}91.42$}\ \ \ $\bm{{\color[rgb]{.5,.5,.5}99.36}}\ \ \$$\color[rgb]{.5,.5,.5}90.46$}\ \ \ $\color[rgb]{.5,.5,.5}84.90$}\ \ \ $\color[rgb]{.5,.5,.5}90.97$}\ \ \ $\color[rgb]{.5,.5,.5}91.42$}\ \ \
CDF $\bm{99.25}\downarrow$ $\bm{91.33}\uparrow$ $86.17\uparrow$ $91.74\uparrow$ $2.12\uparrow 99$$9.10\downarrow$ $\bm{91.31}\uparrow$ $85.48\uparrow$ $91.82\uparrow$ $91.93\uparrow$